OpenStudy (anonymous):

Kai cuts out all the page numbers of a book. He takes each number that has more than one digit and cuts the numbers into separate digits. Given that the book has 999 pages numbered 1 through 999, if he chooses one of these digits at random, what is the probability it is a 7?

OpenStudy (here_to_help15):

Every number occurs the 300 times except 0. (1/10 of 3000, including 000 in this count since it doesn't matter) P(7 or any other non-zero digit for that matter) = 300 / 2889 = 100 / 963 = 10.384 % P(0) = 189 / 2889 = ?